OpenStudy (anonymous):

Consider the function f(x) = 6sin(x2) on the interval 0 ≤ x ≤ 3. (a) Find the exact value in the given interval where an antiderivative, F, reaches its maximum. x = (b) If F(1) = 6, estimate the maximum value attained by F. (Round your answer to three decimal places.) y ≈

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6sin%28x^2%29++domain+0..3 will reach the max when the area above the curve is greatest, i.e. at the point where the curve crosses the \(x\) axis in the picture

OpenStudy (anonymous):

since \(\sin(\pi)=0\) your answer is \(x^2=\pi\) or \(x=\sqrt{\pi}\)

OpenStudy (anonymous):

It said it is wrong

OpenStudy (anonymous):

oh nevermind, it worked now. thank you. would you be able to help me with part b?